# The effect of accretion on scalar superradiant instability

### v1

@@ Appendix A Some useful formulae
$g_{011}(x)\equiv$ $\frac{1}{36}\left[\frac{1}{35}\sqrt{3}k\left(216k^{4}+36k^{2}x^{2}+5x^{4}\right)z^{3}+\frac{2}{7}\sqrt{6}kx^{7}-\frac{2\alpha_{0}x^{9}}{3}\right],$ (89a)
$\begin{split}g_{022}(x)\equiv&-1.11\times 10^{-5}\alpha_{0}x^{13}+k\left(2.15\times 10^{-5}-3.68\times 10^{-4}k\alpha_{0}\right)x^{11}+k^{2}\left(2.97\times 10^{-4}z\alpha_{0}-7.58\times 10^{-6}\right)x^{10}\\ &+k^{3}\left(5.24\times 10^{-4}-5.78\times 10^{-4}k\alpha_{0}\right)x^{9}+k^{4}\left(-2.97\times 10^{-4}z\alpha_{0}-8.45\times 10^{-4}\right)x^{8}+3.64\times 10^{-3}k^{5}x^{7}\\ &+k^{6}\left(7.45\times 10^{-4}-3.06\times 10^{-3}z\alpha_{0}\right)x^{6}+k^{7}z\left(1.35\times 10^{-2}-3.3\times 10^{-2}k\alpha_{0}\right)x^{4}\\ &+k^{9}z\left(1.61\times 10^{-1}-3.96\times 10^{-1}k\alpha_{0}\right)x^{2}+k^{11}z\left(2.91-7.13k\alpha_{0}\right),\end{split}$ (89b)
$\begin{split}g_{033}(x)\equiv&-1.82\times 10^{-9}\alpha_{0}x^{17}+k\left(5.06\times 10^{-9}-1.78\times 10^{-7}k\alpha_{0}\right)x^{15}+k^{2}\left(1.44\times 10^{-7}z\alpha_{0}-1.79\times 10^{-9}\right)x^{14}\\ &+k^{3}\left(3.6\times 10^{-7}-9.85\times 10^{-7}k\alpha_{0}\right)x^{13}+k^{4}\left(3.26\times 10^{-6}z\alpha_{0}-5.85\times 10^{-7}\right)x^{12}\\ &+k^{5}\left(9.49\times 10^{-7}-2.07\times 10^{-5}k\alpha_{0}\right)x^{11}+k^{6}\left(1.37\times 10^{-6}z\alpha_{0}-1.02\times 10^{-5}\right)x^{10}\\ &+k^{7}\left(9.45\times 10^{-5}-4.96\times 10^{-6}k\alpha_{0}\right)x^{9}+k^{8}\left(-4.24\times 10^{-5}z\alpha_{0}-2.16\times 10^{-5}\right)x^{8}\\ &+7.81\times 10^{-5}k^{9}x^{7}+k^{10}\left(2.03\times 10^{-4}-4.36\times 10^{-4}z\alpha_{0}\right)x^{6}+k^{11}z\left(2.26\times 10^{-3}-4.71\times 10^{-3}k\alpha_{0}\right)x^{4}\\ &+k^{13}z\left(2.72\times 10^{-2}-5.65\times 10^{-2}k\alpha_{0}\right)x^{2}+k^{15}z\left(4.89\times 10^{-1}-1.02k\alpha_{0}\right),\end{split}$ (89c)
$\begin{split}g_{044}(x)\equiv&-1.15\times 10^{-13}\alpha_{0}x^{21}+k\left(4.15\times 10^{-13}-2.23\times 10^{-11}k\alpha_{0}\right)x^{19}+k^{2}\left(1.8\times 10^{-11}z\alpha_{0}-1.47\times 10^{-13}\right)x^{18}\\ &+k^{3}\left(5.81\times 10^{-11}-2.85\times 10^{-10}k\alpha_{0}\right)x^{17}+k^{4}\left(1.17\times 10^{-9}z\alpha_{0}-9.44\times 10^{-11}\right)x^{16}\\ &+k^{5}\left(-2.6\times 10^{-9}k\alpha_{0}-2.51\times 10^{-10}\right)x^{15}+k^{6}\left(9.61\times 10^{-9}z\alpha_{0}-4.68\times 10^{-9}\right)x^{14}\\ &+k^{7}\left(1.2\times 10^{-8}-1.02\times 10^{-7}k\alpha_{0}\right)x^{13}+k^{8}\left(4.31\times 10^{-8}z\alpha_{0}-3.89\times 10^{-8}\right)x^{12}\\ &+k^{9}\left(4.62\times 10^{-7}-1.73\times 10^{-7}k\alpha_{0}\right)x^{11}+k^{10}\left(-1.69\times 10^{-7}z\alpha_{0}-2.91\times 10^{-7}\right)x^{10}\\ &+k^{11}\left(1.53\times 10^{-6}-1.62\times 10^{-8}k\alpha_{0}\right)x^{9}+k^{12}\left(7.46\times 10^{-7}-1.97\times 10^{-6}z\alpha_{0}\right)x^{8}\\ &+4.77\times 10^{-7}k^{13}x^{7}+k^{14}\left(1.16\times 10^{-5}-2.03\times 10^{-5}z\alpha_{0}\right)x^{6}+k^{15}z\left(1.26\times 10^{-4}-2.19\times 10^{-4}k\alpha_{0}\right)x^{4}\\ &+k^{17}z\left(1.51\times 10^{-3}-2.62\times 10^{-3}k\alpha_{0}\right)x^{2}+k^{19}z\left(2.72\times 10^{-2}-4.72\times 10^{-2}k\alpha_{0}\right),\end{split}$ (89d)
$\begin{split}g_{055}(x)\equiv&-3.38\times 10^{-18}\alpha_{0}x^{25}+k\left(1.5\times 10^{-17}-1.08\times 10^{-15}k\alpha_{0}\right)x^{23}+kz\left(8.75\times 10^{-16}k\alpha_{0}-5.3\times 10^{-18}\right)x^{22}\\ &+k^{3}\left(3.46\times 10^{-15}-2.53\times 10^{-14}k\alpha_{0}\right)x^{21}+k^{4}\left(1.1\times 10^{-13}z\alpha_{0}-5.62\times 10^{-15}\right)x^{20}\\ &+k^{5}\left(4.84\times 10^{-13}k\alpha_{0}-4.87\times 10^{-14}\right)x^{19}+k^{6}\left(2.32\times 10^{-12}z\alpha_{0}-5.37\times 10^{-13}\right)x^{18}\\ &+k^{7}\left(-1.31\times 10^{-11}k\alpha_{0}-4.27\times 10^{-12}\right)x^{17}+k^{8}\left(1.79\times 10^{-11}z\alpha_{0}-9.76\times 10^{-12}\right)x^{16}\\ &+k^{9}\left(6.77\times 10^{-11}-1.63\times 10^{-10}k\alpha_{0}\right)x^{15}+k^{9}z\left(1.32\times 10^{-10}k\alpha_{0}-1.04\times 10^{-10}\right)x^{14}\\ &+k^{11}\left(8.35\times 10^{-10}-8.03\times 10^{-10}k\alpha_{0}\right)x^{13}+k^{12}\left(-1.41\times 10^{-10}z\alpha_{0}-8.63\times 10^{-10}\right)x^{12}\\ &+k^{13}\left(6.29\times 10^{-9}-4.9\times 10^{-10}k\alpha_{0}\right)x^{11}+k^{14}\left(-4.01\times 10^{-9}z\alpha_{0}-1.15\times 10^{-10}\right)x^{10}\\ &+k^{15}\left(7.07\times 10^{-9}-2.45\times 10^{-11}k\alpha_{0}\right)x^{9}+k^{16}\left(2.61\times 10^{-8}-4.07\times 10^{-8}z\alpha_{0}\right)x^{8}+1.16\times 10^{-9}k^{17}x^{7}\\ &+k^{18}\left(2.82\times 10^{-7}-4.18\times 10^{-7}z\alpha_{0}\right)x^{6}+k^{19}z\left(3.04\times 10^{-6}-4.52\times 10^{-6}k\alpha_{0}\right)x^{4}\\ &+k^{21}z\left(3.65\times 10^{-5}-5.42\times 10^{-5}k\alpha_{0}\right)x^{2}+k^{23}z\left(6.57\times 10^{-4}-9.76\times 10^{-4}k\alpha_{0}\right),\end{split}$ (89e)
$\begin{split}g_{066}(x)\equiv&-5.24\times 10^{-23}\alpha_{0}x^{29}+k\left(2.76\times 10^{-22}-2.5\times 10^{-20}k\alpha_{0}\right)x^{27}+kz\left(2.02\times 10^{-20}k\alpha_{0}-9.75\times 10^{-23}\right)x^{26}\\ &+k^{3}\left(9.47\times 10^{-20}-9.3\times 10^{-19}k\alpha_{0}\right)x^{25}+k^{4}\left(4.16\times 10^{-18}z\alpha_{0}-1.54\times 10^{-19}\right)x^{24}\\ &+k^{5}\left(5.65\times 10^{-17}k\alpha_{0}-2.53\times 10^{-18}\right)x^{23}+k^{6}\left(1.63\times 10^{-16}z\alpha_{0}-2.39\times 10^{-17}\right)x^{22}\\ &+k^{7}\left(5.08\times 10^{-16}k\alpha_{0}-5.38\times 10^{-16}\right)x^{21}+k^{8}\left(1.93\times 10^{-15}z\alpha_{0}-7.64\times 10^{-16}\right)x^{20}\\ &+k^{9}\left(-1.65\times 10^{-14}k\alpha_{0}-3.51\times 10^{-15}\right)x^{19}+k^{10}\left(2.5\times 10^{-14}z\alpha_{0}-9.49\times 10^{-15}\right)x^{18}\\ &+k^{11}\left(7.75\times 10^{-14}-1.57\times 10^{-13}k\alpha_{0}\right)x^{17}+k^{12}\left(1.56\times 10^{-13}z\alpha_{0}-1.69\times 10^{-13}\right)x^{16}\\ &+k^{13}\left(1.03\times 10^{-12}-1.19\times 10^{-12}k\alpha_{0}\right)x^{15}+k^{14}\left(1.99\times 10^{-13}z\alpha_{0}-1.14\times 10^{-12}\right)x^{14}\\ &+k^{15}\left(9.33\times 10^{-12}-2.\times 10^{-12}k\alpha_{0}\right)x^{13}+k^{16}\left(-3.97\times 10^{-12}z\alpha_{0}-2.95\times 10^{-12}\right)x^{12}\\ &+k^{17}\left(2.36\times 10^{-11}-6.26\times 10^{-13}k\alpha_{0}\right)x^{11}+k^{18}\left(2.64\times 10^{-11}-4.35\times 10^{-11}z\alpha_{0}\right)x^{10}\\ &+k^{19}\left(1.34\times 10^{-11}-1.94\times 10^{-14}k\alpha_{0}\right)x^{9}+k^{20}\left(3.35\times 10^{-10}-4.36\times 10^{-10}z\alpha_{0}\right)x^{8}\\ &+1.34\times 10^{-12}k^{21}x^{7}+k^{22}\left(3.47\times 10^{-9}-4.48\times 10^{-9}z\alpha_{0}\right)x^{6}+k^{23}z\left(3.74\times 10^{-8}-4.84\times 10^{-8}k\alpha_{0}\right)x^{4}\\ &+k^{25}z\left(4.49\times 10^{-7}-5.81\times 10^{-7}k\alpha_{0}\right)x^{2}+k^{27}z\left(8.08\times 10^{-6}-1.05\times 10^{-5}k\alpha_{0}\right),\end{split}$ (89f)
$\begin{split}g_{077}(x)\equiv&-4.69\times 10^{-28}\alpha_{0}x^{33}+k\left(2.85\times 10^{-27}-3.13\times 10^{-25}k\alpha_{0}\right)x^{31}+k^{2}\left(2.53\times 10^{-25}z\alpha_{0}-1.01\times 10^{-27}\right)x^{30}\\ &+k^{3}\left(1.36\times 10^{-24}-1.69\times 10^{-23}k\alpha_{0}\right)x^{29}+k^{4}\left(7.67\times 10^{-23}z\alpha_{0}-2.22\times 10^{-24}\right)x^{28}\\ &+k^{5}\left(1.94\times 10^{-21}k\alpha_{0}-5.76\times 10^{-23}\right)x^{27}+k^{6}\left(4.78\times 10^{-21}z\alpha_{0}-5.07\times 10^{-22}\right)x^{26}\\ &+k^{7}\left(7.37\times 10^{-20}k\alpha_{0}-2.09\times 10^{-20}\right)x^{25}+k^{8}\left(8.34\times 10^{-20}z\alpha_{0}-2.52\times 10^{-20}\right)x^{24}\\ &+k^{9}\left(-2.36\times 10^{-20}k\alpha_{0}-5.43\times 10^{-19}\right)x^{23}+k^{10}\left(1.18\times 10^{-18}z\alpha_{0}-3.57\times 10^{-19}\right)x^{22}\\ &+k^{11}\left(-7.24\times 10^{-18}k\alpha_{0}-7.36\times 10^{-19}\right)x^{21}+k^{12}\left(1.85\times 10^{-17}z\alpha_{0}-7.52\times 10^{-18}\right)x^{20}\\ &+k^{13}\left(2.94\times 10^{-17}-1.15\times 10^{-16}k\alpha_{0}\right)x^{19}+k^{14}\left(1.12\times 10^{-16}z\alpha_{0}-1.33\times 10^{-16}\right)x^{18}\\ &+k^{15}\left(8.8\times 10^{-16}-8.31\times 10^{-16}k\alpha_{0}\right)x^{17}+k^{15}z\left(3.39\times 10^{-16}k\alpha_{0}-9.62\times 10^{-16}\right)x^{16}\\ &+k^{17}\left(7.05\times 10^{-15}-2.63\times 10^{-15}k\alpha_{0}\right)x^{15}+k^{18}\left(-1.84\times 10^{-15}z\alpha_{0}-3.78\times 10^{-15}\right)x^{14}\\ &+k^{19}\left(2.92\times 10^{-14}-2.19\times 10^{-15}k\alpha_{0}\right)x^{13}+k^{20}\left(1.06\times 10^{-14}-2.69\times 10^{-14}z\alpha_{0}\right)x^{12}\\ &+k^{21}\left(3.64\times 10^{-14}-4.18\times 10^{-16}k\alpha_{0}\right)x^{11}+k^{21}z\left(2.25\times 10^{-13}-2.68\times 10^{-13}k\alpha_{0}\right)x^{10}\\ &+k^{23}\left(1.24\times 10^{-14}-8.81\times 10^{-18}k\alpha_{0}\right)x^{9}+k^{23}z\left(2.34\times 10^{-12}-2.68\times 10^{-12}k\alpha_{0}\right)x^{8}\\ &+8.42\times 10^{-16}k^{25}x^{7}+k^{26}\left(2.4\times 10^{-11}-2.75\times 10^{-11}z\alpha_{0}\right)x^{6}+k^{27}z\left(2.6\times 10^{-10}-2.97\times 10^{-10}k\alpha_{0}\right)x^{4}\\ &+k^{29}z\left(3.11\times 10^{-9}-3.57\times 10^{-9}k\alpha_{0}\right)x^{2}+k^{31}z\left(5.61\times 10^{-8}-6.42\times 10^{-8}k\alpha_{0}\right),\end{split}$ (89g)

### v2

@@ Appendix A Some useful formulae
$g_{011}(x)\equiv$ $\frac{1}{36}\left[\frac{1}{35}\sqrt{3}k\left(216k^{4}+36k^{2}x^{2}+5x^{4}\right)z^{3}+\frac{2}{7}\sqrt{6}kx^{7}-\frac{2\alpha_{0}x^{9}}{3}\right],$ (A.12a)
$\begin{split}g_{022}(x)\equiv&-1.11\times 10^{-5}\alpha_{0}x^{13}+k\left(2.15\times 10^{-5}-3.68\times 10^{-4}k\alpha_{0}\right)x^{11}\\ &+k^{2}\left(2.97\times 10^{-4}z\alpha_{0}-7.58\times 10^{-6}\right)x^{10}+k^{3}\left(5.24\times 10^{-4}-5.78\times 10^{-4}k\alpha_{0}\right)x^{9}\\ &+k^{4}\left(-2.97\times 10^{-4}z\alpha_{0}-8.45\times 10^{-4}\right)x^{8}+3.64\times 10^{-3}k^{5}x^{7}\\ &+k^{6}\left(7.45\times 10^{-4}-3.06\times 10^{-3}z\alpha_{0}\right)x^{6}+k^{7}z\left(1.35\times 10^{-2}-3.3\times 10^{-2}k\alpha_{0}\right)x^{4}\\ &+k^{9}z\left(1.61\times 10^{-1}-3.96\times 10^{-1}k\alpha_{0}\right)x^{2}+k^{11}z\left(2.91-7.13k\alpha_{0}\right),\end{split}$ (A.12b)
$\begin{split}g_{033}(x)\equiv&-1.82\times 10^{-9}\alpha_{0}x^{17}+k\left(5.06\times 10^{-9}-1.78\times 10^{-7}k\alpha_{0}\right)x^{15}\\ &+k^{2}\left(1.44\times 10^{-7}z\alpha_{0}-1.79\times 10^{-9}\right)x^{14}+k^{3}\left(3.6\times 10^{-7}-9.85\times 10^{-7}k\alpha_{0}\right)x^{13}\\ &+k^{4}\left(3.26\times 10^{-6}z\alpha_{0}-5.85\times 10^{-7}\right)x^{12}+k^{5}\left(9.49\times 10^{-7}-2.07\times 10^{-5}k\alpha_{0}\right)x^{11}\\ &+k^{6}\left(1.37\times 10^{-6}z\alpha_{0}-1.02\times 10^{-5}\right)x^{10}+k^{7}\left(9.45\times 10^{-5}-4.96\times 10^{-6}k\alpha_{0}\right)x^{9}\\ &+k^{8}\left(-4.24\times 10^{-5}z\alpha_{0}-2.16\times 10^{-5}\right)x^{8}+7.81\times 10^{-5}k^{9}x^{7}\\ &+k^{10}\left(2.03\times 10^{-4}-4.36\times 10^{-4}z\alpha_{0}\right)x^{6}+k^{11}z\left(2.26\times 10^{-3}-4.71\times 10^{-3}k\alpha_{0}\right)x^{4}\\ &+k^{13}z\left(2.72\times 10^{-2}-5.65\times 10^{-2}k\alpha_{0}\right)x^{2}+k^{15}z\left(4.89\times 10^{-1}-1.02k\alpha_{0}\right),\end{split}$ (A.12c)
$\begin{split}g_{044}(x)\equiv&-1.15\times 10^{-13}\alpha_{0}x^{21}+k\left(4.15\times 10^{-13}-2.23\times 10^{-11}k\alpha_{0}\right)x^{19}\\ &+k^{2}\left(1.8\times 10^{-11}z\alpha_{0}-1.47\times 10^{-13}\right)x^{18}+k^{3}\left(5.81\times 10^{-11}-2.85\times 10^{-10}k\alpha_{0}\right)x^{17}\\ &+k^{4}\left(1.17\times 10^{-9}z\alpha_{0}-9.44\times 10^{-11}\right)x^{16}+k^{5}\left(-2.6\times 10^{-9}k\alpha_{0}-2.51\times 10^{-10}\right)x^{15}\\ &+k^{6}\left(9.61\times 10^{-9}z\alpha_{0}-4.68\times 10^{-9}\right)x^{14}+k^{7}\left(1.2\times 10^{-8}-1.02\times 10^{-7}k\alpha_{0}\right)x^{13}\\ &+k^{8}\left(4.31\times 10^{-8}z\alpha_{0}-3.89\times 10^{-8}\right)x^{12}+k^{9}\left(4.62\times 10^{-7}-1.73\times 10^{-7}k\alpha_{0}\right)x^{11}\\ &+k^{10}\left(-1.69\times 10^{-7}z\alpha_{0}-2.91\times 10^{-7}\right)x^{10}+k^{11}\left(1.53\times 10^{-6}-1.62\times 10^{-8}k\alpha_{0}\right)x^{9}\\ &+k^{12}\left(7.46\times 10^{-7}-1.97\times 10^{-6}z\alpha_{0}\right)x^{8}+4.77\times 10^{-7}k^{13}x^{7}\\ &+k^{14}\left(1.16\times 10^{-5}-2.03\times 10^{-5}z\alpha_{0}\right)x^{6}+k^{15}z\left(1.26\times 10^{-4}-2.19\times 10^{-4}k\alpha_{0}\right)x^{4}\\ &+k^{17}z\left(1.51\times 10^{-3}-2.62\times 10^{-3}k\alpha_{0}\right)x^{2}+k^{19}z\left(2.72\times 10^{-2}-4.72\times 10^{-2}k\alpha_{0}\right),\end{split}$ (A.12d)
$\begin{split}g_{055}(x)\equiv&-3.38\times 10^{-18}\alpha_{0}x^{25}+k\left(1.5\times 10^{-17}-1.08\times 10^{-15}k\alpha_{0}\right)x^{23}\\ &+kz\left(8.75\times 10^{-16}k\alpha_{0}-5.3\times 10^{-18}\right)x^{22}+k^{3}\left(3.46\times 10^{-15}-2.53\times 10^{-14}k\alpha_{0}\right)x^{21}\\ &+k^{4}\left(1.1\times 10^{-13}z\alpha_{0}-5.62\times 10^{-15}\right)x^{20}+k^{5}\left(4.84\times 10^{-13}k\alpha_{0}-4.87\times 10^{-14}\right)x^{19}\\ &+k^{6}\left(2.32\times 10^{-12}z\alpha_{0}-5.37\times 10^{-13}\right)x^{18}+k^{7}\left(-1.31\times 10^{-11}k\alpha_{0}-4.27\times 10^{-12}\right)x^{17}\\ &+k^{8}\left(1.79\times 10^{-11}z\alpha_{0}-9.76\times 10^{-12}\right)x^{16}+k^{9}\left(6.77\times 10^{-11}-1.63\times 10^{-10}k\alpha_{0}\right)x^{15}\\ &+k^{9}z\left(1.32\times 10^{-10}k\alpha_{0}-1.04\times 10^{-10}\right)x^{14}+k^{11}\left(8.35\times 10^{-10}-8.03\times 10^{-10}k\alpha_{0}\right)x^{13}\\ &+k^{12}\left(-1.41\times 10^{-10}z\alpha_{0}-8.63\times 10^{-10}\right)x^{12}+k^{13}\left(6.29\times 10^{-9}-4.9\times 10^{-10}k\alpha_{0}\right)x^{11}\\ &+k^{14}\left(-4.01\times 10^{-9}z\alpha_{0}-1.15\times 10^{-10}\right)x^{10}+k^{15}\left(7.07\times 10^{-9}-2.45\times 10^{-11}k\alpha_{0}\right)x^{9}\\ &+k^{16}\left(2.61\times 10^{-8}-4.07\times 10^{-8}z\alpha_{0}\right)x^{8}+1.16\times 10^{-9}k^{17}x^{7}\\ &+k^{18}\left(2.82\times 10^{-7}-4.18\times 10^{-7}z\alpha_{0}\right)x^{6}+k^{19}z\left(3.04\times 10^{-6}-4.52\times 10^{-6}k\alpha_{0}\right)x^{4}\\ &+k^{21}z\left(3.65\times 10^{-5}-5.42\times 10^{-5}k\alpha_{0}\right)x^{2}+k^{23}z\left(6.57\times 10^{-4}-9.76\times 10^{-4}k\alpha_{0}\right),\end{split}$ (A.12e)
$\begin{split}g_{066}(x)\equiv&-5.24\times 10^{-23}\alpha_{0}x^{29}+k\left(2.76\times 10^{-22}-2.5\times 10^{-20}k\alpha_{0}\right)x^{27}\\ &+kz\left(2.02\times 10^{-20}k\alpha_{0}-9.75\times 10^{-23}\right)x^{26}+k^{3}\left(9.47\times 10^{-20}-9.3\times 10^{-19}k\alpha_{0}\right)x^{25}\\ &+k^{4}\left(4.16\times 10^{-18}z\alpha_{0}-1.54\times 10^{-19}\right)x^{24}+k^{5}\left(5.65\times 10^{-17}k\alpha_{0}-2.53\times 10^{-18}\right)x^{23}\\ &+k^{6}\left(1.63\times 10^{-16}z\alpha_{0}-2.39\times 10^{-17}\right)x^{22}+k^{7}\left(5.08\times 10^{-16}k\alpha_{0}-5.38\times 10^{-16}\right)x^{21}\\ &+k^{8}\left(1.93\times 10^{-15}z\alpha_{0}-7.64\times 10^{-16}\right)x^{20}+k^{9}\left(-1.65\times 10^{-14}k\alpha_{0}-3.51\times 10^{-15}\right)x^{19}\\ &+k^{10}\left(2.5\times 10^{-14}z\alpha_{0}-9.49\times 10^{-15}\right)x^{18}+k^{11}\left(7.75\times 10^{-14}-1.57\times 10^{-13}k\alpha_{0}\right)x^{17}\\ &+k^{12}\left(1.56\times 10^{-13}z\alpha_{0}-1.69\times 10^{-13}\right)x^{16}+k^{13}\left(1.03\times 10^{-12}-1.19\times 10^{-12}k\alpha_{0}\right)x^{15}\\ &+k^{14}\left(1.99\times 10^{-13}z\alpha_{0}-1.14\times 10^{-12}\right)x^{14}+k^{15}\left(9.33\times 10^{-12}-2.\times 10^{-12}k\alpha_{0}\right)x^{13}\\ &+k^{16}\left(-3.97\times 10^{-12}z\alpha_{0}-2.95\times 10^{-12}\right)x^{12}+k^{17}\left(2.36\times 10^{-11}-6.26\times 10^{-13}k\alpha_{0}\right)x^{11}\\ &+k^{18}\left(2.64\times 10^{-11}-4.35\times 10^{-11}z\alpha_{0}\right)x^{10}+k^{19}\left(1.34\times 10^{-11}-1.94\times 10^{-14}k\alpha_{0}\right)x^{9}\\ &+k^{20}\left(3.35\times 10^{-10}-4.36\times 10^{-10}z\alpha_{0}\right)x^{8}+1.34\times 10^{-12}k^{21}x^{7}\\ &+k^{22}\left(3.47\times 10^{-9}-4.48\times 10^{-9}z\alpha_{0}\right)x^{6}+k^{23}z\left(3.74\times 10^{-8}-4.84\times 10^{-8}k\alpha_{0}\right)x^{4}\\ &+k^{25}z\left(4.49\times 10^{-7}-5.81\times 10^{-7}k\alpha_{0}\right)x^{2}+k^{27}z\left(8.08\times 10^{-6}-1.05\times 10^{-5}k\alpha_{0}\right),\end{split}$ (A.12f)
$\begin{split}g_{077}(x)\equiv&-4.69\times 10^{-28}\alpha_{0}x^{33}+k\left(2.85\times 10^{-27}-3.13\times 10^{-25}k\alpha_{0}\right)x^{31}\\ &+k^{2}\left(2.53\times 10^{-25}z\alpha_{0}-1.01\times 10^{-27}\right)x^{30}+k^{3}\left(1.36\times 10^{-24}-1.69\times 10^{-23}k\alpha_{0}\right)x^{29}\\ &+k^{4}\left(7.67\times 10^{-23}z\alpha_{0}-2.22\times 10^{-24}\right)x^{28}+k^{5}\left(1.94\times 10^{-21}k\alpha_{0}-5.76\times 10^{-23}\right)x^{27}\\ &+k^{6}\left(4.78\times 10^{-21}z\alpha_{0}-5.07\times 10^{-22}\right)x^{26}+k^{7}\left(7.37\times 10^{-20}k\alpha_{0}-2.09\times 10^{-20}\right)x^{25}\\ &+k^{8}\left(8.34\times 10^{-20}z\alpha_{0}-2.52\times 10^{-20}\right)x^{24}+k^{9}\left(-2.36\times 10^{-20}k\alpha_{0}-5.43\times 10^{-19}\right)x^{23}\\ &+k^{10}\left(1.18\times 10^{-18}z\alpha_{0}-3.57\times 10^{-19}\right)x^{22}+k^{11}\left(-7.24\times 10^{-18}k\alpha_{0}-7.36\times 10^{-19}\right)x^{21}\\ &+k^{12}\left(1.85\times 10^{-17}z\alpha_{0}-7.52\times 10^{-18}\right)x^{20}+k^{13}\left(2.94\times 10^{-17}-1.15\times 10^{-16}k\alpha_{0}\right)x^{19}\\ &+k^{14}\left(1.12\times 10^{-16}z\alpha_{0}-1.33\times 10^{-16}\right)x^{18}+k^{15}\left(8.8\times 10^{-16}-8.31\times 10^{-16}k\alpha_{0}\right)x^{17}\\ &+k^{15}z\left(3.39\times 10^{-16}k\alpha_{0}-9.62\times 10^{-16}\right)x^{16}+k^{17}\left(7.05\times 10^{-15}-2.63\times 10^{-15}k\alpha_{0}\right)x^{15}\\ &+k^{18}\left(-1.84\times 10^{-15}z\alpha_{0}-3.78\times 10^{-15}\right)x^{14}=+k^{19}\left(2.92\times 10^{-14}-2.19\times 10^{-15}k\alpha_{0}\right)x^{13}\\ &+k^{20}\left(1.06\times 10^{-14}-2.69\times 10^{-14}z\alpha_{0}\right)x^{12}+k^{21}\left(3.64\times 10^{-14}-4.18\times 10^{-16}k\alpha_{0}\right)x^{11}\\ &+k^{21}z\left(2.25\times 10^{-13}-2.68\times 10^{-13}k\alpha_{0}\right)x^{10}+k^{23}\left(1.24\times 10^{-14}-8.81\times 10^{-18}k\alpha_{0}\right)x^{9}\\ &+k^{23}z\left(2.34\times 10^{-12}-2.68\times 10^{-12}k\alpha_{0}\right)x^{8}+8.42\times 10^{-16}k^{25}x^{7}\\ &+k^{26}\left(2.4\times 10^{-11}-2.75\times 10^{-11}z\alpha_{0}\right)x^{6}+k^{27}z\left(2.6\times 10^{-10}-2.97\times 10^{-10}k\alpha_{0}\right)x^{4}\\ &+k^{29}z\left(3.11\times 10^{-9}-3.57\times 10^{-9}k\alpha_{0}\right)x^{2}+k^{31}z\left(5.61\times 10^{-8}-6.42\times 10^{-8}k\alpha_{0}\right),\end{split}$ (A.12g)